OpenStudy (anonymous):

If a snowball melts so that its surface area decreases at a rate of 0.2cm^2/min, find the rate at which the diameter decreases when the diameter is 13cm

OpenStudy (amistre64):

Whats the equation for the surface area of a sphere?

OpenStudy (bahrom7893):

Sa = 4*pi*r^2

OpenStudy (bahrom7893):

Sa = 4*pi*(d/2)^2 = pi*d^2

OpenStudy (bahrom7893):

dSa/dt = 2pi*d*(dd/dt) 0.2 = 2pi*13*(dd/dt)

OpenStudy (bahrom7893):

dd/dt = 0.2/(26pi)

OpenStudy (bahrom7893):

sorry jinnie i gotta go, got a class early in the morning tomorrow

OpenStudy (anonymous):

thanks bahrom, take care